OpenStudy (twizttiez):

Complete the equation for the linear function whose graph contains the points (2, 5) and (3, 0). y − = (x − 2)

OpenStudy (twizttiez):

I think its y-5=3(x-2)?

OpenStudy (danjs):

calculate the slope from those 2 points, and use the point (2,5) in y - y1 = m(x-x1)

OpenStudy (triciaal):

you can start by finding the slope

OpenStudy (danjs):

You have the point (x1,y1) = (2,5) right the slope is the difference in the 2 y's over the diference in the 2 x's

OpenStudy (danjs):

\[slope = \frac{ 0-5 }{ 3-2 }\]

OpenStudy (twizttiez):

-5?

OpenStudy (danjs):

yea

OpenStudy (twizttiez):

So y-(-5)=3(x-2)

OpenStudy (danjs):

Given: y − ? = ? * (x − 2) where did you get the 3 from

OpenStudy (twizttiez):

Complete the equation for the linear function whose graph contains the points (2, 5) and (3, 0). lol *facepalm* i messed up huh? Is this it So y-(-5)=2(x-2)

OpenStudy (danjs):

point (x1,y1) = (2,5) slope m = -5 ----------------- Point-slope formula y - y1 = m*(x-x1)

OpenStudy (danjs):

y - 5 = -5*(x-2)

OpenStudy (twizttiez):

:| I'm dumb huh?

OpenStudy (danjs):

hah, no you are learning

OpenStudy (twizttiez):

Lol i have 2 more thank god!

OpenStudy (danjs):

k